OpenStudy (anonymous):

A football stadium earned $60,000 in sold-out football ticket sales. Upper level tickets cost $25 per seat and lower level tickets cost $100 per seat. The number of upper level seats is four times the number of lower level seats. How many seats are in the upper level?

OpenStudy (bahrom7893):

25x + 100y = 60,000 x = 4y

OpenStudy (bahrom7893):

x - upper level y - lower level

OpenStudy (bahrom7893):

x = 4y, so: 25*4y + 100y = 60,000 100y + 100y = 60,000 200y = 60,000 y = 300 (lower level)

OpenStudy (bahrom7893):

upper level: x = 4*y = 4*300 = 1200

OpenStudy (anonymous):

Ohhh Thanks so much :))

OpenStudy (bahrom7893):

np :))

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

1 YEAR AGO!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!